# Exponential separations between classical and quantum learners

### v1

@@ 3.1 A learning separation based on a worst-case to average-case reduction
3.2 A learning separation based on obfuscation
The cube root concept class was first studied in [KV94b], and the fact that this concept class exhibits a $\mathsf{CC/QC}$ separation was first observed in [SG04], albeit using different terminology. We note that there exist many similar concept classes based on public-key cryptosystems such as the RSA cryptosystem that exhibit $\mathsf{CC/QC}$ separations (see [KV94a]). Recall that for $\mathsf{CC/QC}$ separations the hypothesis class has to be classically evaluatable, so the role of the quantum computer is only to identify which hypothesis is close to the concept that is generating the examples.
Definition 15 (Cube root concept class [KV94b]).
Fix an $n$-bit integer $N=pq$ [^3] Throughout the paper, the integer $N$ is known to the learner beforehand but $p$ and $q$ are not., where $p$ and $q$ are two $\lfloor n/2\rfloor$-bit primes such that $\gcd\big{(}3,(p-1)(q-1)\big{)}=1$. We define the cube root concept class as $\mathcal{C}_{n}^{\mathrm{DCR}}=\{c_{i}\}_{i\in[n]}$, with
$c_{i}(x)=\mathrm{bin}(f_{N}^{-1}(x),i),$
@@ 3.2 A learning separation based on obfuscation
where $\mathrm{bin}(y,i)$ denotes the $i$th bit of the binary representation of $y$, and the function $f_{N}^{-1}$ is the inverse of $f_{N}(x)=x^{3}\mod N$ defined on $\mathbb{Z}^{*}_{N}$ (i.e., the multiplicative group of integers modulo $N$).
Remark.
By requiring that $p$ and $q$ satisfy $\gcd\big{(}3,(p-1)(q-1)\big{)}=1$, we ensure that $f_{N}^{-1}$ exists.
@@
$(x,c_{i}(x))=\left(y^{3},\mathrm{bin}(y,i)\right),$ (11)
where $y\in\mathbb{Z}_{N}^{*}$ is the unique element such that $x\equiv y^{3}\mod N$. Secondly, note that $f_{N}(x)=x^{3}\mod N$ is a bijection from $\mathbb{Z}_{N}^{*}$ to itself, which implies that sampling $x\in\mathbb{Z}_{N}^{*}$ uniformly at random is equivalent to sampling $y\in\mathbb{Z}_{N}^{*}$ uniformly at random and computing $x=y^{3}\mod N$. By combining this observation with Eq. (11), one finds that one can efficiently generate examples of the cube root concept $c_{i}$ under the uniform distribution over $\mathbb{Z}_{N}^{*}$ by first sampling $y\in\mathbb{Z}_{N}^{*}$ uniformly at random, and then computing $y^{3}\mod N$ together with the $i$th bit of the binary representation of $y$.
The weakest hardness assumption that one can leverage to obtain classical non-learnability is the Discrete Cube Root Assumption (DCRA), which states that computing $f_{N}^{-1}$ is classically intractable (i.e., not in $\mathsf{BPP}$). Namely, in [ACGS88, GMT82] it is shown that computing the least-significant bit of $f_{N}^{-1}$ on any $\frac{1}{2}+\frac{1}{\mathrm{poly}(n)}$ fraction of inputs is at least as hard as computing entire binary representation of $f_{N}^{-1}(x)$ on all inputs. Using the terminology of Section 2.2, if one assumes that computing $f_{N}^{-1}$ is classically intractable (i.e., not in $\mathsf{BPP}$), then the concepts $c_{n}\in\mathcal{C}^{\mathrm{DCR}}_{n}$ lie outside of $\mathsf{HeurBPP}$. In conclusion, analogous to the discrete logarithm concept class, to obtain the classical non-learnability it is sufficient to assume that computing $f_{N}^{-1}$ is classically intractable.
We remark that also in this case one could obtain a similar learning separation for the singleton concept class $\mathcal{C}^{\prime}_{n}=\{c_{n}\}$ for the concept $c_{n}\in\mathcal{C}^{\mathrm{DCR}}_{n}$. This singleton concept class is quantum learnable without requiring the example oracle (i.e. without requiring any data), and one could thus argue that it is not a genuine learning problem anymore (i.e., similar to Observation 1).
However, there is a significant distinction between the learning separations for the discrete logarithm concept class and the cube root concept class that is worth highlighting: the latter is quantumly learnable using a classically evaluatable hypothesis class. To see why this is the case, it is important to note that $f_{N}^{-1}$ is of the form
@@
for some $d^{*}$ that only depends on $N$ [^4] In cryptographic terms, $d^{*}$ is the private decryption key corresponding to the public encryption key $e=3$ and public modulus $N$ in the RSA cryptosystem.. The function $f_{N}^{-1}$ is a type of “trap-door function” in that if one is also given $d^{*}$, then computing $f_{N}^{-1}$ suddenly becomes classically tractable. In other words, there exist polynomially-sized Boolean circuits which evaluate this function, whereas for the discrete logarithm we do not know whether such circuits exist. In this example we thus see the relevance of how the concepts are specified. The specifications “$f^{-1}_{N}$ where $f(x)=x^{3}$” and “$f_{N}^{-1}=x^{d^{*}}$” refer to the same functions, yet computing them is in one case classically tractable, and in the other case it is classically intractable (under the DCRA). The ideas of concealing (easy) functions in difficult descriptions is reminiscent of the term “obfuscation” in computer science, and we will use this term in this context as well. Specifically, we say that the specification “$f^{-1}_{N}$ where $f(x)=x^{3}$” is an obfuscation of the specification “$f_{N}^{-1}=x^{d^{*}}$”. Using the terminology of Section 2.2, this establishes that the problem of evaluating the concepts actually lies inside $\mathsf{P/poly}$ (where the advice string – i.e., $d^{*}$ – is used to “de-obfuscate” the function).
@@
Another feature of the cube root concept class which warrants a comment is that even though computing $d^{*}$ does not require access to the example oracle (recall that $N$ is known beforehand), we still have to learn the bit of $x^{d^{*}}$ that is generating the examples, which does require access to the example oracle (i.e., it requires data). We summarize the results regarding the separation of the cube root concept class in the following theorem.
@@
$L_{\mathrm{DCR}}=\big{(}\{\mathcal{C}_{n}^{\mathrm{DCR}}\}_{n\in\mathbb{N}},\{\mathcal{D}^{U}_{n}\}_{n\in\mathbb{N}})$ exhibits a $\mathsf{CC/QC}$ separation, where $\mathcal{D}^{U}_{n}$ denotes the uniform distribution over $\mathbb{Z}_{N}^{*}$.
3.3 A learning separation with efficiently evaluatable concepts
@@
In this section we establish a learning separation (contingent on a plausible though relatively unexplored hardness assumption) where the concepts do not just admit polynomial-sized Boolean circuits, but are also given in a representation which is efficiently evaluatable on a classical computer. For this concept class, the hardness of learning then cannot stem from the hardness of evaluating the concepts, and it thus lies in identifying which specific concept is generating the examples. To the best of our knowledge, no such separation was given in the literature before. The concept class that satisfies all of the above is the modular exponentiation concept class defined as follows.
Definition 16 (Modular exponentiation concept class).
Let $N=pq$ be an $n$-bit $2^{c}$-integer as defined in Definition 17 with $\gcd\big{(}3,(p-1)(q-1)\big{)}=1$. We define the modular exponentiation concept class as
$\mathcal{C}_{n}^{\mathrm{modexp}}=\Big{\{}\text{ }c_{d}\text{ }\enskip\Big{|}\enskip d=1,\dots,(p-1)(q-1)\text{ and }\gcd(d,(p-1)(q-1)=1\Big{\}},$
@@
$c_{d}:\mathbb{Z}_{N}^{*}\rightarrow\mathbb{Z}_{N}^{*},\quad c_{d}(x)=x^{d}\mod N,$ (14)
and $\mathbb{Z}_{N}^{*}$ denotes the multiplicative group of integers modulo $N$.
@@
An $n$-bit integer $N=pq$ is a $2^{c}$-integer if $p$ and $q$ are two $\lfloor n/2\rfloor$-bit primes such that:
@@
We show that the above concept class is not classically learnable under the assumption that computing $f_{N}^{-1}$ is classically intractable (i.e., not in $\mathsf{BPP}$) when we restrict $N$ to be a $2^{c}$ integer. We will refer to this assumption as the $2^{c}$-discrete cube root assumption ($2^{c}$-DCRA). First, note that the modular exponentiation concept class contains the cube root function $f_{N}^{-1}$ discussed in the previous section (though this time it is not “obfuscated”). Moreover, using the construction from the previous section we can efficiently generate examples $(y,f_{N}^{-1}(y))$, for $y\in\mathbb{Z}_{N}^{*}$ uniformly at random. If we put these examples into an efficient classical learning algorithm for the modular exponentiation concept class, it would with high probability identify a classically efficiently evaluatable hypothesis that agrees with $f_{N}^{-1}$ on a $1-\frac{1}{\mathrm{poly}(n)}$ fraction of inputs. Analogous to the previous section, by the worst-case to average-case reduction of [ACGS88, GMT82] this directly violates the $2^{c}$-DCRA.
We note that by imposing that $N$ is a $2^{c}$-integer might cause the $2^{c}$-DCRA to no longer hold, since there could be an efficient classical algorithm for these specific $2^{c}$-integer moduli. However, since $2^{c}$-integers are generally not considered to be unsecure or “weak” moduli for the RSA cryptosystem, and since recently factored RSA numbers [^5] https://en.wikipedia.org/wiki/RSA_numbers are all essentially $2^{c}$-integers, it is plausible that the $2^{c}$-DCRA still holds (see Appendix B.1 for more details).
To show that the modular exponentiation concept class is quantumly learnable, we use a combination of the quantum algorithm for order-finding and the quantum algorithm for the discrete logarithm [Sho99]. The key observation is that an example $(x,x^{d}\text{ mod }N)$ specifies a congruence relation $d\equiv a\text{ mod }r$, where $r$ denotes the multiplicative order of $x\in\mathbb{Z}^{*}_{N}$, and $a$ denotes the discrete logarithm of $x^{d}$ in the subgroup generated by $x$ (i.e., the smallest positive integer $\ell$ such that $x^{\ell}\equiv x^{d}\mod N$). Next, using the fact that $N$ is a $2^{c}$-number, we show that a polynomial number of these congruences suffices to recover $d$ with high probability. We summarize the learning separation of the modular exponentiation concepts in Theorem 3, and defer the proof to Appendix B.
@@
If the $2^{c}$-DCRA holds, then the learning problem $L_{\mathrm{modexp}}=\big{(}\{\mathcal{C}_{n}^{\mathrm{modexp}}\}_{n\in\mathbb{N}},\{\mathcal{D}^{U}_{n}\}_{n\in\mathbb{N}})$ exhibits a $\mathsf{CC/QC}$ separation, where $\mathcal{D}^{U}_{n}$ denotes the uniform distribution over $\mathbb{Z}_{N}^{*}$.
In conclusion, the modular exponentiation concept class exhibits a $\mathsf{CC/QC}$ separation (assuming the $2^{c}$-DCRA hold), where the concepts are classically efficiently evaluatable. Since the concepts are classically efficiently evaluatable, one could argue that the classical hardness lies in identifying rather than evaluating a hypothesis that is close to the concept generating the examples. We remark that for the modular exponentation concept class, it is not possible to restrict the concept class and obtain a similar learning separation where a quantum learner does not require any data (i.e., similar to Observation 1). In fact, since the concepts are efficiently evaluatable, any polynomially-sized subset of concepts is classically learnable since one can simply do a brute-force search to find the concept that best matches the data.

### v2

@@ 3.1 Learning separation based on heuristic hardness: case of discrete logarithm
3.2 Learning separation based on obfuscation: the case of discrete cube root
In this section, we study learning separations for concept classes based on the discrete cube root, as first studied in [KV94b, SG04]. In short, in the discrete cube root problem, one is given some moduli $N$ and some input $x\in\mathbb{Z}_{N}^{*}$, and one is tasked with finding $y\in\mathbb{Z}_{N}^{*}$ such that $y^{3}\equiv x\mod N$. In contrast to the $\mathsf{CC/QQ}$ separations based on the discrete logarithm discussed in the previous section, we will show that the concept classes based on the discrete cube root exhibit a $\mathsf{CC/QC}$ separation. In particular, one can construct a classically efficient hypothesis class for the concept classes based on the discrete cube root. Nonetheless, we will show that one still needs a quantum learning algorithm to select the correct hypothesis from the classically efficient hypothesis class.
The authors of [KV94a] introduce a concept class based on the discrete cube root and show that it not efficiently classically learnable. Additionally, in [SG04] the authors argue that this concept class is efficiently learnable by a quantum learner, and that they thus exhibit a learning separation. However, upon closer inspection it turns out that their proof relies on being able to allow the input distribution to depend on the concept you are trying to learn. In essence, changing the input distribution is used to reveal the moduli $N$ that is being used for a specific concept. Note that this does not fit within the definition of the PAC learning framework in Section 2, where the input distribution has to be the same for all concepts in the concept class. One way to make the definition from [KV94a] comply with our definition of the PAC learning framework without changing the underlying functions, is to fix the moduli for every instance size.
Definition 17 (Cube root concept class).
[…]
$c_{j}(x)=\mathrm{bin}(f_{N_{i}}^{-1}(x),j),$
@@ 3.2 Learning separation based on obfuscation: the case of discrete cube root
where $\mathrm{bin}(y,j)$ denotes the $j$th bit of the binary representation of $y$, and the function $f_{N}^{-1}$ is the inverse of $f_{N}(x)=x^{3}\mod N$ defined on $\mathbb{Z}^{*}_{N}$ (i.e., the multiplicative group of integers modulo $N$).
Remark (Efficient data generation).
@@
$(x,c_{j}(x))=\left(y^{3},\mathrm{bin}(y,i)\right),$ (11)
where $y\in\mathbb{Z}_{N}^{*}$ is the unique element such that $x\equiv y^{3}\mod N$. Secondly, note that $f_{N}(x)=x^{3}\mod N$ is a bijection from $\mathbb{Z}_{N}^{*}$ to itself, which implies that sampling $x\in\mathbb{Z}_{N}^{*}$ uniformly at random is equivalent to sampling $y\in\mathbb{Z}_{N}^{*}$ uniformly at random and computing $x=y^{3}\mod N$. By combining this observation with Eq. (11), one finds that one can efficiently generate examples of the cube root concept $c_{j}$ under the uniform distribution over $\mathbb{Z}_{N}^{*}$ by first sampling $y\in\mathbb{Z}_{N}^{*}$ uniformly at random, and then computing $y^{3}\mod N$ together with the $j$th bit of the binary representation of $y$.
Classically efficient hypotheses
As already mentioned, there is a significant distinction between the learning separations for the discrete logarithm concept class and the cube root concept class that is worth highlighting: the latter is quantumly learnable using a classically evaluatable hypothesis class. To see why this is the case, it is important to note that $f_{N}^{-1}$ is of the form
@@
for some $d^{*}$ that only depends on $N$ [^6] In cryptographic terms, $d^{*}$ is the private decryption key corresponding to the public encryption key $e=3$ and public modulus $N$ in the RSA cryptosystem.. The function $f_{N}^{-1}$ is a type of “trap-door function” in that if one is also given $d^{*}$, then computing $f_{N}^{-1}$ suddenly becomes classically tractable. In other words, there exist polynomially-sized Boolean circuits which evaluate this function, whereas for the discrete logarithm we do not know whether such circuits exist. In this example we thus see the relevance of how the concepts are specified. The specifications “$f^{-1}_{N}$ where $f(x)=x^{3}$” and “$f_{N}^{-1}=x^{d^{*}}$” refer to the same functions, yet computing them is in one case classically tractable, and in the other case it is classically intractable (under the DCRA). The ideas of concealing (easy) functions in difficult descriptions is reminiscent of the term “obfuscation” in computer science, and we will use this term in this context as well. Specifically, we say that the specification ”$f^{-1}_{N}$ where $f(x)=x^{3}$” is an obfuscation of the specification ”$f_{N}^{-1}=x^{d^{*}}$” in the sense that both represent the same function, but one is not only much harder to understand, but in this case also harder to evaluate. Using the terminology of Section 2.2, this establishes that the problem of evaluating the concepts actually lies inside $\mathsf{P/poly}$ (where the advice string – i.e., $d^{*}$ – is used to “de-obfuscate” the function).
@@
Another feature of the cube root concept class which warrants a comment is that even though computing $d^{*}$ does not require access to the example oracle (recall that $N$ is known beforehand), we still have to learn the bit of $x^{d^{*}}$ that is generating the examples, which does require access to the example oracle (i.e., it requires data).
Hardness assumptions
Much like in the case of the discrete logarithm based separations discussed in Section 3.1, while [SG04] suggests a separation for their concept class by leveraging the canonical hardness assumption of the discrete cube root as outlined in [KV94b], we note that we were unable to obtain a learning separation for the concept class in Definition 17 based on the hardness assumption in [KV94b]by following the proof ideas in [SG04]. Nonetheless, we will discuss two ways of obtaining a learning separation for the modified concept class. First, we propose introducing a new and stronger hardness assumption that adequately supports a proof of separation. Next, we will explore an alternative approach to achieve a learning separation by modifying the underlying functions, allowing us to leverage the canonical hardness of the discrete cube root outlined in [KV94b].
As discussed above, one approach to achieve a $\mathsf{CC/QC}$ separation for the modified concept class defined in 17 involves relying on a stronger hardness assumption. Specifically, this assumption (which we denote as $\mathsf{DCR}$-fixed) is that there exists a sequence of moduli $\{N_{i}\}_{i\in\mathbb{N}}$ for which the discrete cube root is classically intractable. We summarize the results regarding the separation of the modified cube root concept class in the following theorem.
@@
$L_{\mathrm{DCR}}=\big{(}\{\mathcal{C}_{i}^{\mathrm{DCR}}\}_{i\in\mathbb{N}},\{\mathcal{D}^{U}_{i}\}_{i\in\mathbb{N}})$ exhibits a $\mathsf{CC/QC}$ separation assuming the intractability of the discrete cube root for the fixed sequence of moduli $\{N_{i}\}_{i\in\mathbb{N}}$, where $\mathcal{D}^{U}_{i}$ denotes the uniform distribution over $\mathbb{Z}_{N_{i}}^{*}$.
Remark (Heuristic hardness).
A worst-case hardness assumption is sufficient for classical non-learnability, as the discrete cube root admits a worst-to-average case reduction [ACGS88, GMT82]. Note that this worst-to-average case reduction holds for both the canonical hardness assumption as outlined in [KV94b] as well as our newly introduced and weaker $\mathsf{DCR}$-fixed assumption.
However, we again note that the $\mathsf{DCR}$-fixed assumption is stronger than the typical one concerning the discrete cube root, as outlined in [KV94b]. In particular, the hardness assumption outlined in [KV94b] states that there does not exist a polynomial-time classical algorithm that can compute the discrete cube root for most moduli, not that there does not exist a polynomial-time algorithm for any fixed sequence of moduli. This prompts the question of whether learning separations can also be achieved by leveraging the canonical hardness of the discrete cube root outlined in [KV94b]. We show that this is indeed possible and that one can relax the requirement for the existence of sequences of moduli for which the discrete cube root is classically intractable. Specifically, one can define a different yet closely related concept class, that uses a different input space and underlying set of functions, and show that it exhibits a $\mathsf{CC/QC}$ separation assuming the canonical hardness of the discrete logarithm outlined in [KV94b].
Definition 18.
[…]
$c_{(N,i)}(b,j,x)=\begin{cases}\mathrm{bin}(f^{-1}_{N}(x),i)\quad&\text{if }b=0\\ \mathrm{bin}(N,j)\quad&\text{if }b=1.\end{cases}$
for $b\in\{0,1\}$, $j\in[n]=\{1,\dots,n\}$ and $x\in\{0,1\}^{n}$.
This new concept class essentially involves a mechanism for leaking information about the moduli used by the concept in the data. We defer a formal proof of the theorem below to Appendix C.1.
Theorem 5.
$\hat{L}_{\mathrm{DCR}}=\left(\{\hat{\mathcal{C}}^{\mathrm{DCR}}_{n}\}_{n\in\mathbb{N}},\{\mathcal{D}^{U}_{n}\}_{n\in\mathbb{N}}\right)$ exhibits a $\mathsf{CC/QC}$ separation assuming the intractability of the discrete cube root as outlined in [KV94b], where $\mathcal{D}_{n}^{U}$ denotes the uniform distribution over $\mathcal{X}_{n}=\{0,1\}\times[n]\times\{0,1\}^{n}$.
Singleton concept class separations
Another question we would like to address is whether separations are possible for singleton concept classes. As discussed in Section 2, this is crucial in determining whether a problem is a genuine learning problem or merely a computational problem in disguise. Firstly, we observe that a learning separation for the singleton concept class $\mathcal{C}_{i}=\{c_{i}\}$, where $c_{i}\in\mathcal{C}^{\mathrm{DCR}}_{i}$ corresponds to the least significant bit, can be achieved if one assumes our new, yet stronger $\mathsf{DCR}$-fixed hardness assumptions. However, one might wonder whether it is also possible to get a separation for a singleton concept class assuming the canonical hardness assumption of [KV94b]. It turns out that this is indeed the case, though it requires one to define a different yet closely related singleton concept class that utilizes a similar construction to Definition 18.
Definition 19.
We define the concept class $\hat{\mathcal{C}}^{\mathrm{DCR}}_{n}=\{c_{n}\}$, where
$c_{n}(N,x)=\mathrm{bin}(f^{-1}_{N}(x),n)\quad\text{for $N,x\in\{0,1\}^{n}$.}$
Theorem 6.
[…]
A formal proof of the above theorem can be found in Appendix C.2. We note that in our construction the learning separation for a singleton concept class assuming the the canonical hardness assumption of [BM84] comes at the expense of having to change the type of learning separation that is achieved from a $\mathsf{CC/QC}$ separation to a $\mathsf{CC/QQ}$. Therefore, we notice again tradeoff where stronger assumptions are required for $\mathsf{CC/QC}$ separations for binary singleton concept classes, while the difficulty assumption can be relaxed for $\mathsf{CC/QQ}$ separations for binary singleton concept classes
3.3 Learning separation with efficiently evaluatable concepts
@@
In this section we establish a learning separation (contingent on a plausible though relatively unexplored hardness assumption) where the concepts do not just admit polynomial-sized Boolean circuits, but are also given in a representation which is efficiently evaluatable on a classical computer. For this concept class, the hardness of learning them cannot stem from the hardness of evaluating the concepts, and it thus lies in identifying which specific concept is generating the examples. To the best of our knowledge, no such separation was given in the literature before. The concept class that satisfies all of the above is the modular exponentiation concept class defined as follows.
Definition 20.
We define the modular exponentiation concept class as
[…]
@@
$c_{(N,d)}(b,x)=\begin{cases}x^{d}\mod N&\text{if }b=0\\ N&\text{if }b=1.\end{cases}$
for $b\in\{0,1\}$ and $x\in\{0,1\}^{n}$.
@@
An $n$-bit integer $N=pq$ is a $2^{c}$-integer if $p$ and $q$ are two $\lfloor n/2\rfloor$-bit primes such that $\gcd\big{(}3,(p-1)(q-1)\big{)}=1$ and:
@@
We summarize the learning separation of the modular exponentiation concepts in Theorem 7, and defer the proof to Appendix D, and we discuss various aspects of this learning separation below.
@@
$L_{\mathrm{modexp}}=\left(\{\mathcal{C}^{\mathrm{modexp}}_{n}\}_{n\in\mathbb{N}},\{\mathcal{D}^{U}_{n}\}_{n\in\mathbb{N}}\right)$ exhibits a $\mathsf{CC/QC}$ separation assuming the intractability of the discrete cube root as stated in [KV94b] when restricted to $2^{c}$-integer moduli [^7] Here we mean the intractability assumption that states that no classical algorithm can efficiently compute the discrete cube root $f_{N}^{-1}(x)$ for a $\frac{1}{2}+\frac{1}{\mathrm{poly}(n)}$ fraction of all $2^{c}$-integer moduli $N$ and inputs $x\in\mathbb{Z}_{N}^{*}$., where $\mathcal{D}_{n}^{U}$ denotes the uniform distribution over $\mathcal{X}_{n}=\{0,1\}\times\{0,1\}^{n}$.
Quantum learnability
To show that the modular exponentiation concept class is quantumly learnable, we use a combination of the quantum algorithm for order-finding and the quantum algorithm for the discrete logarithm [Sho99]. The key observation is that an example $(x,x^{d}\text{ mod }N)$ specifies a congruence relation $d\equiv a\text{ mod }r$, where $r$ denotes the multiplicative order of $x\in\mathbb{Z}^{*}_{N}$, and $a$ denotes the discrete logarithm of $x^{d}$ in the subgroup generated by $x$ (i.e., the smallest positive integer $\ell$ such that $x^{\ell}\equiv x^{d}\mod N$). Next, using the fact that $N$ is a $2^{c}$-number, we show that a polynomial number of these congruences suffices to recover $d$ with high probability.
Hardness assumption
We show that the above concept class is not classically learnable assuming the intractability of the discrete cube root as stated in [KV94b] restricted to $2^{c}$-integer moduli77footnotemark: 7. We will refer to this assumption as the $2^{c}$-discrete cube root assumption ($2^{c}$-DCRA). To see this, note that the modular exponentiation concept class contains the cube root function $f_{N}^{-1}$ discussed in Section 3.2 (though this time it is not “obfuscated”). Moreover, using the construction also outlined in Section 3.2 we can efficiently generate examples $(y,f_{N}^{-1}(y))$, for $y\in\mathbb{Z}_{N}^{*}$ uniformly at random. If we put these examples into an efficient classical learning algorithm for the modular exponentiation concept class, it would with high probability identify a classically efficiently evaluatable hypothesis that agrees with $f_{N}^{-1}$ on a $1-\frac{1}{\mathrm{poly}(n)}$ fraction of inputs. Similar to Section 3.2, by the worst-case to average-case reduction of [ACGS88, GMT82] this then violates the $2^{c}$-DCRA.
We note that by imposing that $N$ is a $2^{c}$-integer might cause the $2^{c}$-DCRA to no longer hold, since there could be an efficient classical algorithm for these specific $2^{c}$-integer moduli. However, since $2^{c}$-integers are generally not considered to be unsecure or “weak” moduli for the RSA cryptosystem, and since recently factored RSA numbers [^8] https://en.wikipedia.org/wiki/RSA_numbers are all essentially $2^{c}$-integers, it is plausible that the DCRA still holds when restricted to $2^{c}$-integers (see Appendix D.1 for more details).
Conclusion
The modular exponentiation concept class thus exhibits a $\mathsf{CC/QC}$ separation (assuming the $2^{c}$-DCRA hold), where the concepts are classically efficiently evaluatable. Since the concepts are classically efficiently evaluatable, one could argue that the classical hardness of learning lies in identifying rather than evaluating a hypothesis that is close to the concept generating the examples. We remark that for the modular exponentation concept class, it is not possible to restrict the concept class and obtain a similar learning separation where a quantum learner does not require any data (i.e., similar to Observation 1). In fact, since the concepts are efficiently evaluatable classically, any polynomially-sized subset of concepts is classically learnable since a classical learning algorithm can do a brute-force search to find the concept that best matches the data.